Elise:
Forty ounces of a 35% gold alloy are mixed with 62 ounces of a 20% alloy. Find the concentration of the resulting gold alloy.
1
Expert's answer
2012-07-17T08:15:38-0400
So in forty ounces there are 40*0,35 ounces of pure gold. In 62 ounces there are 62*0,2& ounces of pure gold. So resulting concentration is (40*0,35+62*0,2)/(40+62)*100%=(14+12,4)/102*100%=25,9%